OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)? Answer y – 5 = –1(x – 2) y + 8 = –1(x – 1) y + 5 = 1(x + 2) y – 8 = 1(x – 2)

OpenStudy (anonymous):

first one do the slope formula

OpenStudy (anonymous):

THE Y1-Y2=M (X1-X2)?

OpenStudy (hoblos):

no it is y -Y = m(x-X) [ X & Y are coordinates of one of the 2 points] m=(Y2 - Y1)/(X2 - X1)

OpenStudy (mertsj):

It is the first one.